# Conformal prediction for Deep Classifier via Label Ranking

### v1

@@ Abstract
Conformal prediction is a statistical framework that generates prediction sets containing ground-truth labels with a desired coverage guarantee. The predicted probabilities produced by machine learning models are generally miscalibrated, leading to large prediction sets in conformal prediction. In this paper, we empirically and theoretically show that disregarding the probabilities’ value will mitigate the undesirable effect of miscalibrated probability values. Then, we propose a novel algorithm named Sorted Adaptive prediction sets (SAPS), which discards all the probability values except for the maximum softmax probability. The key idea behind SAPS is to minimize the dependence of the non-conformity score on the probability values while retaining the uncertainty information. In this manner, SAPS can produce sets of small size and communicate instance-wise uncertainty. Theoretically, we provide a finite-sample coverage guarantee of SAPS and show that the expected value of set size from SAPS is always smaller than APS. Extensive experiments validate that SAPS not only lessens the prediction sets but also broadly enhances the conditional coverage rate and adaptation of prediction sets.
@@ 1 Introduction
Machine learning is being deployed in many high-stakes tasks, such as autonomous driving (Bojarski et al., 2016), medical diagnostics (Caruana et al., 2015) and financial decision-making. The trust and safety in these applications are critical, as any erroneous prediction can be costly and dangerous. To assess the reliability of predictions, a popular solution is to quantify the model uncertainty, such as confidence calibration (Guo et al., 2017), MC-Dropout (Gal & Ghahramani, 2016), and Bayesian neural network (Smith, 2013; Blundell et al., 2015). However, these methods lack theoretical guarantees of model performance. This gives rise to the importance of Conformal Prediction (CP) (Vovk et al., 2005; Shafer & Vovk, 2008; Balasubramanian et al., 2014; Angelopoulos & Bates, 2021), which yields prediction sets containing ground-truth labels with a desired coverage guarantee.
@@
In the literature, CP algorithms design non-conformity scores to quantify the degree of deviation between a new instance and the training data, determining the size of the final prediction sets. A higher non-conformity score is associated with a larger prediction set or region, indicating a lower level of confidence in the prediction. For example, Adaptive Prediction Sets (APS) (Romano et al., 2020) calculates the score by accumulating the sorted softmax values in descending order. However, the softmax probabilities typically exhibit a long-tailed distribution, allowing for easy inclusion of those tail classes in the prediction sets. To alleviate this issue, Regularized Adaptive Prediction Sets (RAPS) (Angelopoulos et al., 2021b) exclude unlikely classes by appending a penalty to classes beyond some specified threshold. The non-conformity score of RAPS still involves in unreliable softmax probabilities, leading to suboptimal performance in conformal prediction. This motivates our question: does the probability value play a critical role in conformal prediction?
@@
In this work, we show that the value of softmax probability might be redundant information for constructing the non-conformity score in conformal prediction. We provide an empirical analysis by removing the exact value of softmax probability while preserving the relative rankings of labels. The results indicate that APS using label ranking yields much smaller prediction sets than APS using the softmax outputs, at the same coverage rate. Theoretically, we show that, by removing the probability value, the size of prediction sets generated by APS is consistent with model prediction accuracy. In other words, a model with higher accuracy can produce smaller prediction sets, using APS without access to the probability value. The details of the analysis are presented in Subsection 3.1.
@@
Inspired by the analysis, our key idea is to minimize the dependence of the non-conformity score on the probability values, while retaining the uncertainty information. Specifically, we propose Sorted Adaptive prediction sets (dubbed SAPS), which discards all the probability values except for the maximum softmax probability in the construction of non-conformity score. This can be achieved by replacing the non-maximum probability values with a constant, after sorting in descending order. In effect, SAPS can not only produce sets of small size but also communicate instance-wise uncertainty. Theoretically, we show that the expected value of set size from SAPS is always smaller than APS, using a well-calibrated model.
@@
To verify the effectiveness of our method, we conduct thorough empirical evaluations on common benchmarks, including CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), and ImageNet (Deng et al., 2009). The results demonstrate that SAPS achieves superior performance over the compared methods, including APS and RAPS. For example, our approach reduces the average size of prediction sets from 20.95 to 2.98 – only $\frac{1}{7}$ of the prediction set size from APS. Compared to RAPS, we show that SAPS not only produces a higher conditional coverage rate but also exhibits better adaptability to the instance difficulty.
@@
    We empirically show that the probability value is not necessary in APS. Specifically, APS without probability value generates smaller prediction sets than vanilla APS. Moreover, we theoretically show that APS without probability value can provide stable prediction sets, in which the set size is consistent with the prediction accuracy of models.
@@
    We propose a novel non-conformity score–SAPS that minimizes the dependency on probability value while retaining the uncertainty information. We provide theoretical analyses to show the marginal coverage properties of SAPS and the advantage over APS.
@@
    Extensive experimental results demonstrate the effectiveness of our proposed method. We show that SAPS not only lessens the prediction sets but also broadly enhances the conditional coverage rate and adaptation of prediction sets.
@@
    We provide analyses to improve our understanding of the proposed method. In particular, we contrast with a special variant of RAPS and demonstrate the advantages of our method. We also investigate the effect of calibration on our method.
@@ 2 PRELIMINARIES
In this work, we consider the multi-class classification task with $K$ classes. Let $\mathcal{X}\subset\mathbb{R}^{d}$ be the input space and $\mathcal{Y}:=\{1,\dots,K\}$ be the label space. We use $\hat{\pi}:\mathcal{X}\rightarrow\mathbb{R}^{K}$ to denote the pre-trained neural network that is used to predict the label of a test instance. Let $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ denote a random data pair satisfying a joint data distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$. Ideally, $\hat{\pi}_{y}(\bm{x})$ can be used to approximate the conditional probability of class $i$ given feature $\bm{x}$, i.e., $\mathbb{P}[Y=y|X=\bm{x}]$. Then, the model prediction in classification tasks is generally made as: $\hat{y}=\underset{y\in\mathcal{Y}}{\arg\max}~{}\hat{\pi}_{y}(\bm{x})$.
@@
Conformal prediction. To provide a formal guarantee for the model performance, conformal prediction (Vovk et al., 2005) is designed to produce prediction sets containing ground-truth labels with a desired probability. Instead of predicting one-hot labels from the model outputs, the goal of conformal prediction is to construct a set-valued mapping $\mathcal{C}:\mathcal{X}\rightarrow 2^{\mathcal{Y}}$, which satisfies the marginal coverage:
$\mathbb{P}(Y\in\mathcal{C}_{1-\alpha}(X))\geq 1-\alpha,$ (1)
@@
where $\alpha\in(0,1)$ denotes the desired error rate and $\mathcal{C}_{1-\alpha}(X)$ is a subset of $\mathcal{Y}$. Particularly, a smaller value of $\alpha$ will enlarge the predictions set, i.e.,
$\alpha_{1}>\alpha_{2}\Longrightarrow\mathcal{C}_{1-\alpha_{1}}(X)\subseteq\mathcal{C}_{1-\alpha_{2}}(X)$ (2)
@@
Before deployment, conformal prediction begins with a calibration step, using a calibration set $\mathcal{D}_{cal}={(\bm{x}_{i},y_{i})}_{i=1}^{n}$. The data of the calibration set is also i.i.d. drawn from the distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$. Specifically, we calculate a non-conformity score $s_{i}=S(\bm{x}_{i},y_{i})$ for each example $(\bm{x}_{i},y_{i})$ in the calibration set, where $s_{i}$ measures the degree of deviation between the given example and the training data. The $1-\alpha$ quantile of the non-conformity scores $\{s_{i}\}_{i=1}^{n}$ is then determined as a threshold $\tau$. Formally, the value of $\tau$ can be obtained as shown below:
$\tau=\inf\{s:\frac{|\{s_{i}:s_{i}\leq s,i=1,\dots,n\}|}{n}\geq\frac{\lceil(n+1)(1-\alpha)\rceil}{n}\}$
@@
During testing, we calculate the non-conformity score for each class given a new instance $\bm{x}_{n+1}$. Then, the corresponding prediction set $\mathcal{C}(\bm{x}_{n+1})$ comprises possible labels whose non-conformity score $S(\bm{x}_{n+1},y)$ falls within the threshold $\tau$:
$\mathcal{C}_{1-\alpha}(\bm{x}_{n+1})=\{y:S(\bm{x}_{n+1},y)\leq\tau,y\in\mathcal{Y}\}$ (3)
@@
With a lower value of $\tau$, the model tends to produce a smaller prediction set, indicating a higher level of confidence in the prediction. Conversely, the increase of $\tau$ will enlarge the size of the prediction set, suggesting greater uncertainty of the prediction. In this manner, conformal prediction can be used to estimate the uncertainty or reliability of the model’s predictions.
[…]
@@
Figure 1: (a) Sorted softmax probabilities of an example from ImageNet in descending order. (b) Set size for APS on various models. We use ”w/ value” and ”w/o value” to represent the vanilla APS and APS with label ranking, respectively. The numbers in brackets represent the prediction accuracy of the model. The sizes of the prediction sets are small after removing the probability value.
@@
Adaptive prediction sets (APS). In the APS method (Romano et al., 2019), the non-conformity scores are calculated by accumulating softmax probabilities in descending order. Formally, given a data pair $(\bm{x},y)$, the non-conformity score can be computed by:
$S(\bm{x},y,u;\hat{\pi}):=\sum_{i=1}^{o(y,\hat{\pi}(\bm{x}))-1}\hat{\pi}_{(i)}(\bm{x})+u\cdot\hat{\pi}_{(o(y,\hat{\pi}(\bm{x})))}(\bm{x}),$ (4)
@@
where $o(y,\hat{\pi}(\bm{x}))$ denotes the index of $\hat{\pi}_{y}(\bm{x})$ in the sorted softmax probabilities,i.e., $\hat{\pi}_{(1)}(\bm{x})),\dots,\hat{\pi}_{(K)}(\bm{x}))$ , and $u$ is an independent random variable satisfying a uniform distribution on $[0,1]$. Given an oracle model $\pi$ that outputs the true conditional probabilities, the non-conformity score $S(X,Y,U;\pi)$ can be uniformly distributed in $[0,1]$, resulting in $\tau$ being equal to $1-\alpha$ (Einbinder et al., 2022b). With the non-conformity score in Eq. 4, APS achieves a finite-sample marginal coverage guarantee. However, the softmax probabilities $\hat{\pi}(\bm{x})$ typically exhibit a long-tailed distribution, where the tail probabilities with small values can be easily included in the prediction sets. Consequently, APS tends to produce large prediction sets for all inputs, regardless of the instance difficulty. For example, in Figure 1, the long-tail probability distribution results in the non-conformity scores of many classes falling within $\tau$. This motivates our analysis to investigate the role of probability value in conformal prediction.
@@ 3.1 Motivation
To analyze the role of probability values, we perform an ablation study by removing the influence of probability values in Eq. 4. In particular, we replace these probabilities with a constant $\gamma$ (e.g., $\gamma=1$), after sorting in descending order. With the constant $\gamma$, the modified non-conformity score for a data pair $(\bm{x},y)$ with a pre-trained model $\hat{\pi}$ is:
$S(\bm{x},y,u;\hat{\pi}):=\gamma\cdot\left[o(y,\hat{\pi}(\bm{x}))-1+u\right].$ (5)
@@
In the analysis, we fix the constant as 1 for simplification. Then, we conduct experiments on ImageNet (Deng et al., 2009) to compare the new non-conformity score to the vanilla APS. Here, we set the desired error rate as $10\%$, i.e., $\alpha=0.1$. Following previous works (Romano et al., 2019; Angelopoulos et al., 2021b; Ghosh et al., 2023), we first randomly split the test dataset of ImageNet into two subsets: a conformal calibration subset of size 30K and a test subset of size 20K. For network architecture, we use seven models trained on ImageNet, with different levels of prediction performance (see Figure 1). All models are calibrated by the temperature scaling procedure (Guo et al., 2017). Finally, experiments are repeated ten times and the median results are reported.
@@
Probability values are not necessary. Figure 1 presents the results on various models, using APS with/without the probability value. The results indicate that APS solely based on label ranking generates smaller prediction sets than those generated with the vanilla APS, across various models. For example, with the Inception model, removing the probability values reduces the set size of $88.18$ to $6.33$. Using a transformer-based ViT model (Touvron et al., 2021), APS without probability value also obtains a smaller set size. From the comparison, we show that the probability value might be redundant information for non-conformity scores in conformal prediction. We proceed by theoretically analyzing the advantage of removing probability values in APS.
@@
A theoretical interpretation. The empirical results above demonstrate that the probability value is not a critical component of the non-conformity score for conformal prediction. Here, we provide a formal analysis of APS without probability value through the following theorem:
Theorem 1.
@@
Let $A_{r}$ denote the accuracy of the top $r$ predictions on a trained model $\hat{\pi}$. Given a significance level $\alpha$, for any test instance $\bm{x}\sim\mathcal{P}_{\mathcal{X}}$ and an independent random variable $u\sim U[0,1]$, if there exists a number $k$ satisfying $A_{k}\geq 1-\alpha>A_{k-1}$, the size of prediction set generated by APS without probability value can be obtained by
@@
$|\mathcal{C}_{1-\alpha}(\bm{x},u)|=\left\{\begin{aligned} &k,\quad\quad\mathrm{if}\hskip 2.84526ptu<\frac{1-\alpha-A_{k-1}}{A_{k}-A_{k-1}},\\ &k-1,\quad\mathrm{otherwise}.\end{aligned}\right.$ (6)
@@
$\mathbb{E}_{u\sim[0,1]}[\mathcal{C}_{1-\alpha}(\bm{x},u)]=(k-1)+\frac{1-\alpha-A_{k-1}}{A_{k}-A_{k-1}}.$ (7)
The proof of Theorem 1 can be found in Appendix A. As indicated by Eq. 7, the prediction set size generated by APS without probability value is consistent with $k$. In other words, a higher model accuracy will lead to a smaller value of $k$, indicating a smaller prediction sets. This argument is clearly supported by experimental results shown in Figure 1. In particular, we observe that using APS without probability value, models with higher accuracy produce a smaller prediction sets, while the vanilla APS does not exhibit this characteristic. For example, using ResNeXt101, the model achieves higher prediction accuracy than using ResNet152, while producing a larger prediction set. The analysis demonstrates the advantage of removing probability value in APS, via decreasing the sensitivity to tail probabilities.

### v2

@@ Abstract
Conformal prediction is a statistical framework that generates prediction sets containing ground-truth labels with a desired coverage guarantee. The predicted probabilities produced by machine learning models are generally miscalibrated, leading to large prediction sets in conformal prediction. To address this issue, we propose a novel algorithm named Sorted Adaptive Prediction Sets (SAPS), which discards all the probability values except for the maximum softmax probability. The key idea behind SAPS is to minimize the dependence of the non-conformity score on the probability values while retaining the uncertainty information. In this manner, SAPS can produce compact prediction sets and communicate instance-wise uncertainty. Extensive experiments validate that SAPS not only lessens the prediction sets but also broadly enhances the conditional coverage rate of prediction sets.
Machine Learning, ICML
@@ 1 Introduction
Machine learning is being deployed in many high-stakes tasks, such as autonomous driving (Bojarski et al., 2016), medical diagnostics (Caruana et al., 2015), and financial decision-making. The trust and safety in these applications are critical, as any erroneous prediction can be costly and dangerous. To assess the reliability of predictions, a popular solution is to quantify the model uncertainty, such as confidence calibration (Guo et al., 2017), MC-Dropout (Gal & Ghahramani, 2016), and Bayesian neural network (Smith, 2013). However, these methods lack theoretical guarantees of the model performance. This gives rise to the importance of Conformal Prediction (CP) (Vovk et al., 2005; Shafer & Vovk, 2008; Balasubramanian et al., 2014; Angelopoulos & Bates, 2021), which yields prediction sets containing ground-truth labels with a desired coverage guarantee.
@@
To determine the size of the prediction set, CP algorithms generally design non-conformity scores to quantify the deviation degree between a new instance and the training data. A higher non-conformity score is associated with a larger prediction set or region, indicating a lower confidence level in the prediction. For example, Adaptive Prediction Sets (APS) (Romano et al., 2020) calculates the score by accumulating the sorted softmax values in descending order. However, the softmax probabilities typically exhibit a long-tailed distribution, allowing for easy inclusion of those tail classes in the prediction sets. To alleviate this issue, Regularized Adaptive Prediction Sets (RAPS) (Angelopoulos et al., 2021) excludes unlikely classes by appending a penalty to classes beyond a specified threshold. Yet, the non-conformity score of RAPS involves unreliable softmax probabilities, leading to suboptimal performance in conformal prediction. This motivates our question: does the probability value play a critical role in conformal prediction?
@@
In this work, we show that the value of softmax probability might be redundant information for conformal prediction. We provide an empirical analysis by removing the exact value of softmax probability while preserving the relative rankings of labels. The results indicate that APS without probability value yields much more compact prediction sets than APS using the softmax outputs, at the same coverage rate. Theoretically, we show that, by removing the probability value, the size of prediction sets generated by APS is consistent with the model prediction accuracy. In other words, a model with higher accuracy can produce smaller prediction sets, using APS without access to the probability value. The details of the analysis are presented in Section 3.1.
@@
Inspired by the analysis, our key idea is to minimize the dependence of the non-conformity score on the probability values, while retaining the uncertainty information. Specifically, we propose the Sorted Adaptive prediction sets (dubbed SAPS), which discards all the probability values except for the maximum softmax probability in the construction of non-conformity score. After sorting in descending order, this can be achieved by replacing the non-maximum probability values with a constant. In effect, SAPS can not only produce sets of small size but also communicate instance-wise uncertainty. In addition, we show that as the constant approaches infinity, the expected value of set size from SAPS is asymptotically equivalent to that of APS without probability value (see Figure 3(b)).
@@
To verify the effectiveness of our method, we conduct thorough empirical evaluations on common benchmarks, including CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), and ImageNet (Deng et al., 2009). The results demonstrate that SAPS is superior to the compared methods, including APS and RAPS. For example, our approach reduces the average size of prediction sets from 20.95 to 2.98 – only $\frac{1}{7}$ of the prediction set size from APS. Compared to RAPS, we show that SAPS produces a higher conditional coverage rate and exhibits better adaptability to the instance difficulty. Code are publicly available at https://github.com/ml-stat-Sustech/conformal_prediction_via_label_ranking.
@@
    We show that the probability value might be unnecessary for APS. Specifically, APS without probability value generates smaller prediction sets than vanilla APS. Theoretically, we show that this variant of APS can provide stable prediction sets, in which the set size is consistent with the prediction accuracy of models.
@@
    We propose a novel non-conformity score – SAPS that minimizes the dependency on probability value while retaining the uncertainty information.
@@
    Extensive experimental results demonstrate the effectiveness of our proposed method. We show that SAPS lessens the prediction sets and broadly enhances the conditional coverage rate.
@@
    We provide analyses to improve our understanding of the proposed method. In particular, we validate that both retaining the maximum softmax probability and confidence calibration are important for SAPS to generate compact prediction sets (see Section 5).
@@ 2 Preliminaries
In this work, we consider the multi-class classification task with $K$ classes. Let $\mathcal{X}\subset\mathbb{R}^{d}$ be the input space and $\mathcal{Y}:=\{1,\dots,K\}$ be the label space. We use $\hat{\pi}:\mathcal{X}\rightarrow\mathbb{R}^{K}$ to denote the pre-trained neural network that is used to predict the label of a test instance. Let $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ denote a random data pair satisfying a joint distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$. Ideally, $\hat{\pi}_{y}(\boldsymbol{x})$ can be used to approximate the conditional probability of the class $y$ given an image feature $\boldsymbol{x}$, i.e., $\mathbb{P}[Y=y|X=\boldsymbol{x}]$. Then, the model prediction in classification tasks is generally made as: $\hat{y}=\underset{y\in\mathcal{Y}}{\arg\max}~{}\hat{\pi}_{y}(\boldsymbol{x})$.
Conformal prediction.
@@
To provide a formal guarantee for the model performance, conformal prediction (Vovk et al., 2005) is designed to produce prediction sets containing ground-truth labels with a desired probability. Instead of predicting one-hot labels from the model outputs, the goal of conformal prediction is to construct a set-valued mapping $\mathcal{C}:\mathcal{X}\rightarrow 2^{\mathcal{Y}}$, which satisfies the marginal coverage:
$\mathbb{P}(Y\in\mathcal{C}(X))\geq 1-\alpha,$ (1)
@@
where $\alpha\in(0,1)$ denotes the desired error rate and $\mathcal{C}(X)$ is a subset of $\mathcal{Y}$.
@@
Before deployment, conformal prediction begins with a calibration step, using the calibration set $\mathcal{D}_{cal}:=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}$. The samples of the calibration set are i.i.d., having been drawn from the distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$. Specifically, we calculate a non-conformity score $s_{i}=S(\boldsymbol{x}_{i},y_{i})$ for each example $(\boldsymbol{x}_{i},y_{i})$ in the calibration set, where $s_{i}$ measures the degree of deviation between the given example and the training data. The $1-\alpha$ quantile of the non-conformity scores $\{s_{i}\}_{i=1}^{n}$ is then determined as a threshold $\tau$. Formally, the value of $\tau$ can be obtained as shown below:
$\tau=\inf\{s:\frac{|\{i\in\{1,\dots,n\}:s_{i}\leq s\}|}{n}\geq\frac{\lceil(n+1)(1-\alpha)\rceil}{n}\}.$
@@
During testing, we calculate the non-conformity score for each label given a new instance $\boldsymbol{x}_{n+1}$. Then, the corresponding prediction set $\mathcal{C}(\boldsymbol{x}_{n+1})$ comprises possible labels whose non-conformity score $S(\boldsymbol{x}_{n+1},y)$ falls within $\hat{\tau}$:
$\mathcal{C}_{1-\alpha}(\boldsymbol{x}_{n+1};\tau):=\{y\in\mathcal{Y}:S(\boldsymbol{x}_{n+1},y)\leq\tau\}.$ (3)
The above equation exhibits a nesting property of threshold, i.e.,
$\tau_{1}\leq\tau_{2}\Longrightarrow\mathcal{C}_{1-\alpha}(\boldsymbol{x}_{n+1};\tau_{1})\subseteq\mathcal{C}_{1-\alpha}(\boldsymbol{x}_{n+1};\tau_{2}).$ (4)
@@
This property shows that with a lower value of $\tau$, the model tends to produce a smaller prediction set, indicating a higher level of confidence in the prediction. Conversely, the increase of $\tau$ will enlarge the size of the prediction set, suggesting greater uncertainty of the prediction. In this manner, conformal prediction can be used to estimate the uncertainty or reliability of the model’s predictions.
((a))
((b))
@@
Figure 1: (a) Softmax probabilities for an instance from ImageNet are arranged in descending order. (b) Set size on various models. We use “w/ value” and “w/o value” to represent the vanilla APS and APS without probability values, respectively. The numbers in brackets represent the accuracy of the model. The sizes of the prediction sets are decreased after removing the probability value.
Adaptive prediction sets (APS).
@@
In the APS method (Romano et al., 2020), the non-conformity scores are calculated by accumulating the softmax probabilities in descending order. Formally, given a data pair $(\boldsymbol{x},y)$, the non-conformity score can be computed by:
$S_{aps}(\boldsymbol{x},y,u;\hat{\pi}):=\sum\limits_{i=1}^{o(y,\hat{\pi}(\boldsymbol{x}))-1}\hat{\pi}_{(i)}(\boldsymbol{x})+u\cdot\hat{\pi}_{(o(y,\hat{\pi}(\boldsymbol{x})))}(\boldsymbol{x}),$
where $o(y,\hat{\pi}(\boldsymbol{x}))$ denotes the index of $\hat{\pi}_{y}(\boldsymbol{x})$ in the sorted softmax probabilities, i.e., $\hat{\pi}_{(1)}(\boldsymbol{x}),\dots,\hat{\pi}_{(K)}(\boldsymbol{x})$ , and $u$ is an independent random variable satisfying a uniform distribution on $[0,1]$. Similarly, the prediction set of APS with the error rate $\alpha$ for a test instance $\boldsymbol{x}_{n+1}$ is given by
$\mathcal{C}_{1-\alpha}(\boldsymbol{x}_{n+1},u_{n+1};\tau):=\{y\in\mathcal{Y}:S_{aps}(\boldsymbol{x}_{n+1},y,u_{n+1};\hat{\pi})\leq\tau\},$
where $u_{n+1}\sim U[0,1]$ and $\tau$ is defined by Equation 2. In particular, a larger value of $\tau$ will lead to a larger prediction set, representing higher uncertainty. Thus, the prediction set defined by Equation 6 satisfies the nesting property in Equation 4.
In addition, the prediction set $\mathcal{C}_{1-\alpha}(\boldsymbol{x}_{n+1},u_{n+1};\tau)$ with the $\tau$ obtained by Equation 2 has a finite-sample marginal coverage guarantee, as stated in the following theorem.
Theorem 2.1.
(Angelopoulos et al., 2021) Suppose the calibration data $(X_{i},Y_{i},U_{i})_{i=1,\dots,n}$ and a test instance $(X_{n+1},Y_{n+1},U_{n+1})$ are exchangeable. Let the set-valued function $\mathcal{C}_{1-\alpha}(\boldsymbol{x},u;\tau)$ satisfy the nesting property of $\tau$ in Equation 4. For $\tau$ defined in Equation 2, we have the following coverage guarantee:
$P\left(Y_{n+1}\in\mathcal{C}_{1-\alpha}\left(X_{n+1},U_{n+1};\tau\right)\right)\geq 1-\alpha.$
@@
In practice, the assumption of exchangeability in the above Theorem is promising as it is weaker than the i.i.d. assumption (Teng et al., 2023). However, the softmax probabilities $\hat{\pi}(\boldsymbol{x})$ typically exhibit a long-tailed distribution, where the tail probabilities with small values can be easily included in the prediction sets. Consequently, APS tends to produce large prediction sets for all inputs, regardless of the instance difficulty. For example, in Figure 1(a), the long-tail probability distribution results in the non-conformity scores of many classes falling within the threshold $\tau$. This motivates our analysis to investigate the role of probability values in conformal prediction.
@@ 3.1 Motivation
To analyze the role of probability values, we perform an ablation study by removing the influence of probability values in Equation 5. In particular, after arranging softmax probabilities in descending order, we substitute them with a positive constant $\gamma$ (e.g., $\gamma=1$).
Formally, the modified non-conformity score for a data pair $(\boldsymbol{x},y)$ with a pre-trained model $\hat{\pi}$ can be give by:
$S_{cons}(\boldsymbol{x},y,u;\hat{\pi},\gamma):=\gamma\cdot\left[o(y,\hat{\pi}(\boldsymbol{x}))-1+u\right].$ (7)
In what follows, we thoroughly probe the impacts of $\gamma$. We start by giving the following definition.
Lemma 3.1.
(Vovk et al., 2005) Let $\mathcal{Z}$ be a space of examples. Given a significance level, two non-conformity score functions $A$ and $B$ could provide the same prediction sets if $A(z)\leq A(z^{\prime})\Longleftrightarrow B(z)\leq B(z^{\prime})$ for all $z,z^{\prime}\in\mathcal{Z}$.
Lemma 3.1 states that in conformal prediction, different non-conformity score functions are equivalent if they lead to the same non-conformity order for any data examples.
Proposition 3.2.
For all $\gamma_{1},\gamma_{2}>0$, the score functions $S_{cons}(\boldsymbol{x},y,u;\hat{\pi},\gamma_{1})$ and $S_{cons}(\boldsymbol{x},y,u;\hat{\pi},\gamma_{2})$ could provide the same prediction sets in conformal prediction.
@@
The proof of the above proposition is presented in Appendix A. According to Proposition 3.2, we can find that the value of $\gamma$ in $S_{cons}(\boldsymbol{x},y,u;\hat{\pi},\gamma)$ is irrelevant with the prediction sets. Thus, we fix the constant $\gamma$ to 1 for simplification. We conduct experiments on ImageNet (Deng et al., 2009) to compare the new non-conformity score to the vanilla APS. Here, we set the desired error rate as $10\%$, i.e., $\alpha=0.1$. Following previous works (Romano et al., 2019; Angelopoulos et al., 2021; Ghosh et al., 2023), we first randomly split the test dataset of ImageNet into two subsets: a conformal calibration subset of size 30K and a test subset of size 20K. For the network architecture, we use seven models trained on ImageNet, with different levels of prediction performance (see Figure 1(b)). All models are calibrated by the temperature scaling procedure (Guo et al., 2017) on the calibration dataset. Finally, experiments are repeated ten times and the median results are reported.
Probability values may not be necessary.
@@
Figure 1(b) presents the results on various models, using APS with/without the probability value. The results indicate that APS solely based on label ranking generates smaller prediction sets than the vanilla APS, across various models. For example, on the Inception model, removing the probability values reduces the set size from $88.18$ to $6.33$. Similarly, when applying a transformer-based ViT model (Touvron et al., 2021), APS without probability values also obtains a smaller set size. This comparison suggests that the probability values might be redundant for non-conformity scores in conformal prediction. We further explore this by theoretically analyzing the benefits of discarding softmax probability values in APS.
A theoretical interpretation.
@@
The above empirical results demonstrate that the probability value is not a critical component of the non-conformity score for conformal prediction. Here, we provide a formal analysis of APS without probability value through the following theorem:
Theorem 3.3.
@@
Let $A_{r}$ denote the accuracy of the top $r$ predictions for a trained model $\hat{\pi}$ on an infinite calibration set. Given a significance level $\alpha$, there exists an integer $k$ satisfying $A_{k}\geq 1-\alpha>A_{k-1}$. For any test instance $\boldsymbol{x}\sim\mathcal{P}_{\mathcal{X}}$ and an independent random variable $u\sim U[0,1]$, the size of the prediction set $\mathcal{C}_{1-\alpha}(\boldsymbol{x},u)$ generated by APS without probability value can be obtained by
@@
$|\mathcal{C}_{1-\alpha}(\boldsymbol{x},u)|=\left\{\begin{aligned}&k,\quad\quad\mathrm{if}\hskip 2.84526ptu<\frac{1-\alpha-A_{k-1}}{A_{k}-A_{k-1}},\\ &k-1,\quad\mathrm{otherwise}.\end{aligned}\right.$ (8)
@@
$\mathbb{E}_{u\sim[0,1]}[|\mathcal{C}_{1-\alpha}(\boldsymbol{x},u)|]=k-1+\frac{1-\alpha-A_{k-1}}{A_{k}-A_{k-1}}.$ (9)
The proof of Theorem 3.3 can be found in Appendix B. This is consistent with the theoretical results of RAPS, which indicates that the size of prediction sets generated by RAPS with heavy regularization will be at most equal to $k$ (Angelopoulos et al., 2021). As indicated by Equation 9, the prediction set size generated by APS without probability value is consistent with $k$. In other words, a higher model accuracy will lead to a smaller value of $k$, indicating smaller prediction sets. This argument is clearly supported by experimental results shown in Figure 1(b). In particular, we observe that using APS without probability value, models with higher accuracy produce smaller prediction sets, while the vanilla APS does not exhibit this characteristic. For example, the vanilla APS yields larger prediction sets for ResNeXt101 than for ResNet152, despite ResNeXt101 achieving higher predictive accuracy than ResNet152. The analysis demonstrates the advantage of removing probability value in APS, via decreasing the sensitivity to tail probabilities.